OpenStudy (anonymous):

Will give medal and fan A line has a slope of 9/4 and a y intercept of -14. which equations represent the line? Choose ALL answers that are correct A. -9x + 4y = -56 B. 9x + 4y = 56 C. y = 9/4x -14 D. y = 9/4x + 14

OpenStudy (anonymous):

wait did I close this

OpenStudy (anonymous):

C is correct. Let me find the other correct one.

OpenStudy (anonymous):

The other correct one is A.

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

C is in slope intercept form. A is in standard form. You don't need to solve C, but solving A helps you to know it's correct. -9x + 4y = -56. +9x - 9x + 4y = -56 + 9x 4y= 9x - 56 Divide four by both sides. y= 9/4 - 14. And presto. Same as C when in slope intercept form. (:

OpenStudy (anonymous):

You're very welcome.

OpenStudy (anonymous):

Can you help with some more?